The scrap value at the end of its useful life is given by S(n)=C(1-r)^n, Where C is the original cost, n is the useful life of the machine in years, and r is the constant annual percentage of value lost. Find the scrap value of the original machine. Original Cost: $52,000; Life: 11 years; annual rate of value lost:11%
Answer by stanbon(75887)

The scrap value at the end of its useful life is given by S(n)=C(1-r)^n, Where C is the original cost, n is the useful life of the machine in years, and r is the constant annual percentage of value lost.
-----------------------
Find the scrap value of the original machine. Original Cost: $52,000; Life: 11 years; annual rate of value lost:11%
S(11) = 52000*(1-0.11)^11
S(11) = 52000*0.89^11
S(11) = 52000*0.2775
S(11) = $14,430.90
==================
Cheers,
Stan H.
------------------
------